OpenStudy (anonymous):

solve each quadratic equation. Show your work . (2x – 1)(x + 7) = 0

OpenStudy (ssddss02):

what do you think?

OpenStudy (s4sensitiveandshy):

\[a*b=0\] same as a=0 and b =0 so set both parentheses equal to zero 2nd)solve for x

OpenStudy (skypetutor):

(2x – 1)(x + 7) = 0 So, either 2x - 1 = 0 or x + 7 = 0 That gives us, x = 1/2 and x = -7 Hoping this will help you.

OpenStudy (skypetutor):

Let me know if this helps you.

OpenStudy (anonymous):

sorry i was busy, but thanks tho @skypetutor

OpenStudy (skypetutor):

You are most welcome